We are reattempting the database move for Scrooge to bring improved performance to all users. We anticipate the move taking all night into the early morning. As always, service will be restored by 9AM Eastern. We will update this NOC post with any updates along the way.

Update 4:00 AM Eastern: The database copy is still underway. At this point we are planning on letting the DB copy finish throughout the Friday day as we cannot afford to let the DB copy extend past this weekend.

Users on Scrooge will see Empty mailboxes throughout the day, however rest assured that your previous mailbox is still intact. Over the weekend we will converge all user mailboxes.

Clients using Outlook should have their cached profiles throughout the day that will enable them to work like normal. Clients using OWA are recommended to concurrently use LiveArchive to access mail prior to the DB move.

Update 5:10 AM Eastern: We’ve decided not to reset user mailboxes, and instead run of the current drive for one more day. We understand that users will continue to experience delays, however we feel it would be in the best interest of all users to complete the DB move over the weekend instead of doing a converge.

Update 5:28 AM Eastern: We’ve re-enabled the previous DBs. For the time being, public folders will be unavailable on the server to try and save as much reads/writes as possible on the server until the backup restarts tonight.

Friday 9/3/10

Update 11:42 PM Eastern: We are about to restart the database migration. We anticipate the migration taking at least 12 hours. Customers should use LiveArchive if they need access to their live mail.

Sunday 9/5/10

Update 1:25 PM Eastern: The database backup was successfully completed and we began the process of restoring the backup to the new RAID array. We anticipate mailboxes being available by 7PM Eastern tonight and public folders remounted tomorrow afternoon.

Update 5:45 PM Eastern: We’ve had a few inquires from partners on the lack of updates with the work done yesterday and today. While we love keeping this NOC blog updated with fairly live updates, the majority of time in the past two days has been watching databases and users copy between drives. We do appreciate everyones patience as we attempt to finalize all changes by the end of the holiday weekend.

Update 6:57 PM Eastern: The second attempt at the restore is still in progress. The earlier restore slowed down about half way through, so we restarted the server and disconnected the older RAID array. The current restore is looking good and is estimated to complete in 45 minutes. After the restore is complete, we are going to try to mount the database and then move the databases to it’s final location.

Update 7:50 PM Eastern: The restore was completed and we’ve enabled OWA access to Scrooge. All pending mail is being delivered to the user mailboxes and Outlook access will be enabled after the queues are flushed out. We’re going to test a DB move to the array at 10 PM, and  once an estimated time is established we will update this post.

Update 9:27 PM Eastern: Service has been restored and we’ve postponed the test move and we will update this post.